Ekson Energy, Inc. (EKS) is an energy company with a market debt-to-equity ratio of 3. Suppose its current cost of debt capital is 6%, and its cost of equity capital is 14%.

(a) Calculate EKS’s weighted average cost of capital assuming no taxes and asymmetric information, i.e. Modigliani Miller propositions I and II with no taxes are in place.

Part 2:

Suppose now that if EKS issues equity and uses the proceeds to repay its debt and reduce its debt-toequity ratio to 2, it will lower its debt cost of capital to 5.5%.

(b) Calculate the new WACC assuming MM propositions with no taxes hold.

(c) Calculate EKS’s cost of equity capital after this reduction in leverage.

Part 3:

Assume now that EKS issues even more equity and pays off its debt completely.

(d) Calculate EKS’s cost of equity capital now. (Notice that in this situation EKS is unlevered!)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

& VD ( 1x D ENG - 01:18 = 17-03-2020 ... o X GL1163 X for - GF V GG GH GI GJ GK GL GM GN GO GPA D/E = Wd =D/V = We =E/V = 3 3